# Position: We Need An Adaptive Interpretation of Helpful, Honest, and Harmless Principles

### v3

@@ 7 Open Challenges
Additionally, the framework faces implementation challenges–inherent difficulties associated with multi-objective optimization (MOO). Balancing helpfulness, honesty, and harmlessness requires trade-offs that are complicated by normative assumptions embedded in the optimization objectives. The lack of precise quantification methods for the interaction between these principles complicates achieving a theoretically optimal balance, limiting the applicability of the framework in certain high-stakes or ambiguous scenarios. Following up recent studies propose potential methods for MOO alignment (Bai et al., 2022a; Mukherjee et al., 2024), more efforts should be made in addressing this challenge.
@@
Even though the proposed framework faces these challenges and may be judged by some other communities, we believe that offering a structured reference is a necessary step forward. By fostering discussion, guiding practical implementation, and encouraging interdisciplinary collaboration, this framework can serve as a foundation for future refinements.
@@ 8 Conclusion
The Helpful, Honest, and Harmless (HHH) principles are critical for aligning AI models with human values, but their static and ambiguous interpretation falls short to address the complexities of diverse real-world contexts. This position paper advocates for an adaptive approach to the HHH principle, enabling AI systems to contextually balance these dimensions while maintaining their ethical integrity and practical effectiveness. By introducing a flexible prioritization framework, we aim to bridge the gap between theoretical alignment and real-world applicability, ensuring AI systems remain both responsible across various domains.
Impact Statement
The introduction of an adaptive framework for interpreting the HHH principles redefines AI alignment by addressing ambiguities and contextual variability inherent in current interpretations. This work enhances the operational effectiveness of HHH principles by enabling AI systems to dynamically balance these priorities across domains such as healthcare, education, and creative industries. Its significance lies in ensuring AI systems remain safe, trustworthy, and user-centric while addressing complex tasks, particularly in high-risk applications like medical diagnostics. By promoting transparency and ethical alignment, this framework bridges the gap between theoretical principles and practical application, paving the way for AI models to safeguard societal well-being and foster innovation.
Acknowledgement
We would like to express our sincere gratitude to Prof. Xiuying Chen from MBZUAI for her valuable suggestions and insightful feedback on this paper. Her expertise and thoughtful guidance greatly contributed to the improvement of our work. Max Lamparth is partially supported by the Stanford Center for AI Safety, the Center for International Security and Cooperation, and the Stanford Existential Risk Initiative.
Appendix A Model and Benchmark Example
@@ Appendix B Example of Priority Order
Appendix C Example of Helpfulness Definition
Context Definition of Helpfulness
Appendix D Case Study of the 3H principles for developing Chemistry Foundation Models
@@
Benchmarking Standards Context-Specific Benchmarks: Design benchmarks that align with the prioritized dimensions. For instance: 1). Measured by the model’s predictive accuracy and relevance to user tasks (Guo et al., 2023), 2). Evaluated through safety checks on generated molecular outputs. 3). Assessed by comparing generated outputs with verified scientific data (Guo et al., 2024a; Bushuiev et al., 2024). Multi-Dimensional Integration: Use complementary benchmarks to capture performance holistically across different dimensions.
System Transparency and Governance Transparency: In the chemistry domain, models must provide interpretable outputs, highlighting key chemical features influencing predictions and annotating data sources such as PubChem (Kim et al., 2021) or ChEMBL (Mendez et al., 2019). Confidence scores and error margins should accompany predictions to ensure reliability. Governance: It requires strict validation of outputs against experimental data, safeguards against misuse (e.g., toxic compound generation), and compliance with safety standards like EPA(Environmental Protection Agency)) and REACH(Registration, Evaluation, Authorization, and Restriction of Chemicals). Regular audits by chemists and toxicologists are essential to refine predictions and maintain ethical and safety standards.

### v4

@@ 7 Open Challenges
Additionally, the framework faces implementation challenges–inherent difficulties associated with multi-objective optimization (MOO). Balancing helpfulness, honesty, and harmlessness requires trade-offs that are complicated by normative assumptions embedded in the optimization objectives. The lack of precise quantification methods for the interaction between these principles complicates achieving a theoretically optimal balance, limiting the framework’s applicability in certain high-stakes or ambiguous scenarios. Following up recent studies propose potential methods for MOO alignment (Bai et al., 2022a; Mukherjee et al., 2024), more efforts should be made in addressing this challenge.
@@
Notably, the criticism that HHH principles are merely high-level heuristics and lack operational substance neglects two important facts. First, the original intention of HHH was to anchor alignment efforts in human-centered values, offering interpretable and ethically grounded objectives for training and evaluation. While it is true that these principles are polysemantic and context-sensitive-as are any high-level ethical ideals such as autonomy or respect-this does not diminish their value. Rather, it reinforces the necessity of the adaptive framework we propose, which systematically resolves ambiguities through contextual modeling, priority ordering, and risk assessment. Second, asserting that operational guidance has moved beyond HHH (e.g., to model specs or rule-based reward shaping) overlooks the fact that these newer approaches often implicitly instantiate the HHH values. The 30-40 page model specs (OpenAI, 2024a) and deliberative alignment strategies do not replace HHH; they concretize it. Our contribution lies in formalizing this process: instead of discarding HHH for being abstract, we operationalize it adaptively-preserving its normative clarity while enabling contextual application.
Despite these limitations, we believe a structured reference is a necessary step forward. By guiding implementation, fostering discussion, and enabling interdisciplinary collaboration, this framework can serve as a foundation for future refinement and extension.
@@ 8 Conclusion
The Helpful, Honest, and Harmless (HHH) principles are central to aligning AI with human values. This paper proposes an adaptive framework that enables context-sensitive balancing of HHH, ensuring ethical integrity and practical effectiveness across diverse applications.
Appendix A Model, Framework and Benchmark Example
Framework & Principle Trust. Utility
HHH (Askell et al., 2021) ˇ ˇ
TrustLLM (Huang et al., 2024a) ˇ ˚
NIST (NIST, 2024) ˇ ˚
AI Safety and Security (Qi et al., 2024) ˇ ˘
Risk Assessment (Kapoor et al., 2024) ˇ ˘
Transparency (Bommasani et al., 2024) ˇ ˘
Table 2: Typical principles or frameworks for guiding AI model alignment (excluding government laws or acts, e.g., EU AI Act (EU, 2024b) and AI Bill of Rights (House, 2022)). ˇ, ˚, and ˘ indicate different emphasis on trustworthiness or utility.
@@ Appendix B Example of Priority Order
Appendix C Case Study of the 3H principles for developing Chemistry Foundation Models
@@
Alignment Auditing Context-Specific Benchmarks: Design benchmarks that align with the prioritized dimensions. For instance: 1). Measured by the model’s predictive accuracy and relevance to user tasks (Guo et al., 2023), 2). Evaluated through safety checks on generated molecular outputs. 3). Assessed by comparing generated outputs with verified scientific data (Guo et al., 2024a; Bushuiev et al., 2024). Multi-Dimensional Integration: Use complementary benchmarks to capture performance holistically across different dimensions.
Governance Infrastructure Transparency: In the chemistry domain, models must provide interpretable outputs, highlighting key chemical features influencing predictions and annotating data sources such as PubChem (Kim et al., 2021) or ChEMBL (Mendez et al., 2019). Confidence scores and error margins should accompany predictions to ensure reliability. Governance: It requires strict validation of outputs against experimental data, safeguards against misuse (e.g., toxic compound generation), and compliance with safety standards like EPA(Environmental Protection Agency)) and REACH(Registration, Evaluation, Authorization, and Restriction of Chemicals). Regular audits by chemists and toxicologists are essential to refine predictions and maintain ethical and safety standards.